the number of ordered pairs (x,y) of real numbers that atisfy the simultaneous equations
x+y2=x2+y=12  is

Grade:12

1 Answers

Arun
6 years ago
x2– y2
 = x – y
(x – y)(x + y) = x – y
x = y or, x + y = 1
x + x2= 12
x = –4, 3
The required ordered pairs are (–4, –4), (3, 3), & two irrational roots from x2 – x – 11 = 0